OpenStudy (anonymous):

Find the slope of a line perpendicular to a line through the given points. G(4, -2), H(8, 0)

OpenStudy (anonymous):

is it 2?

OpenStudy (zepp):

\(\large \frac{\Delta y}{\Delta x}=\frac{0-(-2)}{8-4}=\frac{2}{4}=\frac{1}{2}\) And you do a magic trick, -2.

OpenStudy (anonymous):

is this one -1?? Find the slope of a line perpendicular to a line through the given points. U(6, 6), V(4, 4)

Parth (parthkohli):

Yep. -1 is correct

Parth (parthkohli):

:D

OpenStudy (anonymous):

:Dthanks!!